### v1

# KVComm: Enabling Efficient LLM Communication through Selective KV Sharing

Abstract
Large Language Models (LLMs) are increasingly deployed in multi-agent systems, where effective inter-model communication is crucial. Existing communication protocols either rely on natural language, incurring high inference costs and information loss, or on hidden states, which suffer from information concentration bias and inefficiency. To address these limitations, we propose KVComm, a novel communication framework that enables efficient communication between LLMs through selective sharing of KV pairs. KVComm leverages the rich information encoded in the KV pairs while avoiding the pitfalls of hidden states. We introduce a KV layer-wise selection strategy based on attention importance scores with a Gaussian prior to identify the most informative KV pairs for communication. Extensive experiments across diverse tasks and model pairs demonstrate that KVComm achieves comparable performance to the upper-bound method, which directly merges inputs to one model without any communication, while transmitting as few as 30% of layers’ KV pairs. Our study highlights the potential of KV pairs as an effective medium for inter-LLM communication, paving the way for scalable and efficient multi-agent systems.
1 Introduction
Large Language Models (LLMs) have catalyzed a paradigm shift from isolated model capabilities towards collaborative multi-agent systems (Guo et al., 2024; Tran et al., 2025). CAMEL (Li et al., 2023), AutoGen (Wu et al., 2024), and ChatDev (Qian et al., 2023) have demonstrated the potential of LLMs to collaborate effectively in multi-agent systems, achieving impressive results in various tasks. These systems leverage the strengths of individual LLMs and enable them to work together to solve complex problems that are beyond the capabilities of a single model (Yang et al., 2024a).
While multi-agent systems have shown great promise, they also introduce new challenges, particularly in the area of inter-agent communication. Effective communication between LLMs is crucial for the success of multi-agent systems. Explicit communication through natural language has been explored in several works, enabling the models to share information (Du et al., 2023), coordinate their actions (Sun et al., 2025), and make collective decisions (Yang et al., 2024b).
However, natural language communication leads to high inference costs due to the need for multiple decoding steps, and may not fully capture the rich information that needs to be shared between LLMs as information is lost in the sampling process (Pham et al., 2023; Ramesh & Li, 2025) that occurs as each new token is produced. To address this limitation, recent works have explored alternative communication protocols that leverage the internal representations of LLMs. CIPHER (Pham et al., 2023) proposed to use the embedding space as the medium of communication between LLMs. Namely, they pass the weighted average of the token embeddings from one LLM to another, facilitating more efficient information exchange. Rather than using the embedding space, AC (Ramesh & Li, 2025) transmits the intermediate activations, specifically the last token’s hidden state. They replace the last token’s hidden state of the receiver’s model ($\mathcal{M}_{r}$) with that of the sender’s model ($\mathcal{M}_{s}$), allowing a more direct transfer of information. While these methods have shown promising results, they still face challenges in terms of communication efficiency and effectiveness. CIPHER (Pham et al., 2023) still requires multiple decoding steps, which can be costly, and AC (Ramesh & Li, 2025) may lead to information loss as only limited activation information is transmitted.
We start with the question: What is the most effective way to communicate between LLMs? We argue that an ideal communication protocol should satisfy the following criteria: ① Effectiveness: It should enable $\mathcal{M}_{r}$ to effectively utilize the information from $\mathcal{M}_{s}$. ② Efficiency: It should minimize the computation needed by $\mathcal{M}_{s}$ and the amount of data transmitted between models. ③ Generality: It should be applicable to a wide range of tasks and model architectures, ensuring its versatility in different scenarios. We choose to use activation information as the medium of communication, as no decoding steps are needed for $\mathcal{M}_{s}$, and $\mathcal{M}_{r}$ can directly utilize the rich information encoded in the activations. We study different types of activation information (i.e., hidden states and KV pairs), and in Section 2.2, we show that hidden states suffer from information concentration bias, where the last token’s hidden state contains most of the information needed for the model’s output. This makes it challenging to design an effective communication protocol using the last token’s hidden state. Furthermore, we find that using all tokens’ hidden states does not guarantee effective communication. A dilemma arises: if the hidden states are taken from the early layers of $\mathcal{M}_{s}$, the computation benefit is limited since the computation cost is similar to concatenating the two inputs; if the hidden states are prepended to the later layers of $\mathcal{M}_{r}$, the performance drops significantly.
Based on these observations, we propose KVComm, a novel communication protocol that enables efficient communication between LLMs through selective sharing of KV pairs. KV pairs are the most representative activation information in each layer, and sharing them does not interact with the hidden states of $\mathcal{M}_{r}$ directly, while $\mathcal{M}_{r}$ can decide how to utilize the information through the attention mechanism. To further improve the efficiency of communication, we propose a selection strategy to choose which (potentially non-contiguous) layers’ KV pairs to share. We formulate hypotheses that (H1) KV pairs from intermediate layers encode transferable semantic knowledge, and (H2) KV pairs from layers exhibiting stronger attention distributions are more effective for communication. These hypotheses are validated by our experiments in Sections 4.3 and 4.5. Based on these hypotheses, we define attention importance scores for each layer based on the average attention weights assigned to the context tokens. We also apply a Gaussian distribution centered at a certain layer as a prior on the attention importance scores. The intuition is that the Gaussian distribution encourages selecting layers around a certain depth, which aligns with hypothesis H1. The general framework is illustrated in Figure 1.
Figure 1: KVComm framework for efficient LLM communication through selective KV sharing.
We evaluate KVComm on a diverse set of tasks with eight model pairs (see Section 4.1), showing that it consistently outperforms existing communication protocols while significantly reducing the data transmitted between models. In summary, our work makes three key contributions:
  • We evaluate different types of activation information for communication between LLMs, and identify the limitations of using hidden states as the medium of communication. We show that the last token’s hidden state suffers from information concentration bias, and point out a dilemma that arises when using all tokens’ hidden states.
  • We propose KVComm, a novel communication protocol that enables efficient communication between LLMs through selective sharing of KV pairs. We design a selection strategy based on attention importance scores and a Gaussian prior to choose which layers’ KV pairs to share. This is the first approach that makes it possible to choose non-contiguous layers of KV. Moreover, we show the feasibility of using a single context/question pair for guiding the selection for a given pair of models, prior to deployment.
  • We conduct extensive experiments on a diverse set of tasks and model pairs, demonstrating that KVComm enables effective and efficient communication between LLMs, achieving comparable performance to the Skyline method, which is the upper-bound and directly merges the inputs without any communication, while reducing the computation costs by 2.5x to 6x. In particular, KVComm enables up to a 3x reduction in communication relative to approaches that transmit the entire set of KV pairs. Moreover, we demonstrate the performance benefits of non-contiguous selection of KV layers. Finally, we demonstrate the increase in performance that KVComm brings even over Skyline on two datasets, further illustrating the need to communicate in a non-strictly textual manner.
2 Problem and Motivation
2.1 Problem Formulation
We formally define the problem of solving a contextual task through the communication of two LLMs: $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$. $\mathcal{M}_{s}$ takes as input a context $C$, and generates the required information $I_{C}$ to be communicated. $\mathcal{M}_{r}$ takes as input the query $Q$ and the information $I_{C}$ from $\mathcal{M}_{s}$, and produces the final output. In this work, we limit the choices of the two LLMs to (1) two instances of the same LLM, and (2) two models that are fine-tuned versions of the same base LLM. The goal is to design an efficient communication protocol that allows $\mathcal{M}_{s}$ to effectively convey the necessary information to $\mathcal{M}_{r}$ while minimizing the amount of data transmitted.
2.2 Why Hidden States Fall Short
When Decoder-Only LLMs infer, the input information flows through the model in the form of activation values, which refer to the intermediate results output by each decoder layer during the forward pass. We refer to the intermediate activation values that are passed between adjacent layers as hidden states. We also consider the KV pairs used in the attention mechanism within each layer as another type of activation information. In this section, we investigate the effectiveness of using hidden states as the medium of communication by studying two questions: How important are hidden states of tokens at different positions in the sequence? (Section 2.2.1) Are hidden states of all tokens effective for communication? (Section 2.2.2)
2.2.1 Token Importance at Different Positions
We begin with a simple experiment examining how token positions affect performance. Using Llama-3.1-8B on MMLU Social Science, we remove or retain the hidden state of only specific tokens at a given layer and measure the performance change. As shown in Figure 3, different tokens vary in importance across layers, with the last token becoming most critical in later layers. This aligns with the intuition that the last token is often the most relevant to the current prediction. Thus, the last token’s hidden state carries the most influential information for both model output and inter-LLM communication. Results on additional datasets and models are provided in Appendix C.
To ensure efficient communication with hidden states built on this observation, two conditions must hold: (1) $\mathcal{M}_{s}$ must send at least the last token’s hidden state, and (2) the communication protocol should preserve $\mathcal{M}_{r}$’s last token state as much as possible. The protocol in Ramesh & Li (2025) either replaces $\mathcal{M}_{r}$’s last token state with that of $\mathcal{M}_{s}$ or averages the two, but both cause information loss in $\mathcal{M}_{r}$’s last token state, harming its performance.
Figure 2: Compared to other token positions, the last token’s hidden state is the most critical, especially in later layers.
Figure 3: Prepending hidden states is not effective unless hidden states are from and to the early layers.
2.2.2 Utilizing All Tokens
Another straightforward approach to ensure the last token’s hidden state is preserved is to prepend all tokens’ hidden states from $\mathcal{M}_{s}$ to $\mathcal{M}_{r}$. The experiments on HotpotQA with Llama-3.1-8B, presented in Figure 3, demonstrate that prepending all tokens’ hidden states from $\mathcal{M}_{s}$ to $\mathcal{M}_{r}$ is effective if the hidden states are taken from the early layers of $\mathcal{M}_{s}$ and prepended to the early layers of $\mathcal{M}_{r}$. Appendix D shows experimental results on other datasets. We find that this method is caught in a dilemma: (1) if the hidden states are taken from the early layers of $\mathcal{M}_{s}$, the computation benefit is limited since it is similar to concatenating the two inputs; (2) if the hidden states are prepended to the later layers of $\mathcal{M}_{r}$, the performance drops significantly.
These findings suggest that while utilizing all tokens’ hidden states can preserve the last token’s information, it does not guarantee effective communication between LLMs.
3 Efficient LLM Communication through Selective KV Sharing
We propose a simple yet effective communication protocol that enables efficient communication between LLMs by selectively sharing KV pairs. This approach addresses the limitations observed in previous methods by ensuring that the most critical information is preserved. Our design satisfies the three criteria outlined below: it enhances effectiveness by allowing $\mathcal{M}_{r}$ to utilize essential context (①), improves efficiency by reducing unnecessary computation and transmission overhead (②), and ensures generality by being applicable across diverse tasks and architectures (③).
3.1 Communication Framework
For a given context $C$ and query $Q$, $\mathcal{M}_{s}$ processes the context $C$ and runs one forward pass (prefill stage) to generate the KV pairs $\{(\mathbf{k}_{s}^{l},\mathbf{v}_{s}^{l})\}$ at each layer $l$, where $l=1,2,\ldots,L$ and $L$ is the total number of layers in $\mathcal{M}_{s}$. We apply a selection strategy to choose a subset of KV pairs $\{(\mathbf{k}_{s}^{l_{i}},\mathbf{v}_{s}^{l_{i}})\}$, where $i=1,2,\ldots,M$ and $M$ is the number of selected layers. The selected KV pairs are then transmitted to $\mathcal{M}_{r}$.
$\mathcal{M}_{r}$ processes the query $Q$ and incorporates the received KV pairs during its forward passes (prefill and decoding stages). Specifically, at each layer $l$ of $\mathcal{M}_{r}$, if $l$ corresponds to a selected layer $l_{i}$ [^1] The layer indices are 1-to-1 matched between $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$ since we only consider the case where the two models are the same or fine-tuned versions of the same base LLM., the KV pairs from $\mathcal{M}_{s}$ are integrated into the attention mechanism. We simply concatenate the KV pairs from $\mathcal{M}_{s}$ with those of $\mathcal{M}_{r}$: $\mathbf{k}_{r}^{l}\leftarrow[\mathbf{k}_{s}^{l_{i}};\mathbf{k}_{r}^{l}]$, and $\mathbf{v}_{r}^{l}\leftarrow[\mathbf{v}_{s}^{l_{i}};\mathbf{v}_{r}^{l}]$. This integration allows $\mathcal{M}_{r}$ to attend to both its own context and the information provided by $\mathcal{M}_{s}$. After processing the query $Q$ with the integrated KV pairs, $\mathcal{M}_{r}$ generates the final output.
3.2 KV Selection Strategies
The communication protocol critically depends on the selection strategy for choosing which KV pairs to transmit from $\mathcal{M}_{s}$ to $\mathcal{M}_{r}$. Not all layers or attention heads contribute equally to encoding task-relevant knowledge. A fundamental question when designing selection strategies is: Which parts of the KV pairs encode the most relevant knowledge for communication?
Formally, given the set of candidate KV pairs $\{(\mathbf{k}_{s}^{l},\mathbf{v}_{s}^{l})\}_{l=1}^{L}$, our goal is to select a subset $\mathcal{S}\subseteq\{1,\ldots,L\}$ such that the receiver’s output retains maximal information from the sender, given a constraint on the number of selected layers $|\mathcal{S}|=M$, which is determined by the desired communication efficiency. This can be formulated as the following optimization problem:
$\max_{\mathcal{S}\subseteq\{1,\ldots,L\},|\mathcal{S}|=M}f(\mathcal{M}_{r}(Q,\{(\mathbf{k}_{s}^{l},\mathbf{v}_{s}^{l})\}_{l\in\mathcal{S}})),$
where $f(\cdot)$ is a performance metric (e.g., accuracy, F1 score), and $\mathcal{M}_{r}(Q,\{(\mathbf{k}_{s}^{l},\mathbf{v}_{s}^{l})\}_{l\in\mathcal{S}})$ denotes the output of the receiver model given the query $Q$ and the selected KV pairs. Since direct computation of this objective is intractable, we instead propose two hypotheses H1 and H2 that serve as priors for designing practical heuristics.
The first hypothesis H1 is that KV pairs from intermediate layers contain the most readily transferable semantic knowledge. Prior analyses (Jawahar et al., 2019; Geva et al., 2020) suggest a hierarchy: early layers capture surface patterns, middle layers encode semantic abstractions, and late layers specialize in task predictions. Thus, intermediate KV pairs should carry the richest generalizable information, making them most effective for communication. Experiment results in Section 4.3 support this hypothesis.
Another hypothesis H2 is that KV pairs from layers exhibiting stronger attention distributions are more effective for communication. Intuitively, if a head consistently allocates high attention mass to the given tokens, its KV cache encodes salient contextual relations that are critical for the model’s reasoning. Attention concentration thus serves as a proxy for the communication value of a KV subset, suggesting that such heads should be prioritized for selection. This hypothesis is also validated by our experiments in Section 4.5.
Our selection strategy is based on these two hypotheses. We first define attention importance scores for each layer, which are calculated as the average attention weights that have been assigned to the context tokens by all heads in that layer during the prefill stage. We then take a Gaussian distribution centered at a certain layer as a prior to select layers with high attention importance scores. The intuition is that the Gaussian prior encourages selecting layers around a certain depth, which aligns with hypothesis H1 that intermediate layers are more likely to contain transferable knowledge.
Mathematically, the attention importance score for each layer $l$ is computed as:
$\hat{S}_{a}^{l}=\frac{1}{HT}\sum_{h=1}^{H}\sum_{t=1}^{T}\sum_{c=1}^{C}a_{h,t,c}^{l},$
where $H$ is the number of attention heads, $T$ is the number of tokens in the query, $C$ is the number of context tokens, and $a_{h,t,c}^{l}$ is the attention weight assigned by head $h$ at layer $l$ from token $t$ to context token $c$. $\hat{S}_{a}^{l}$ is then normalized to the range $[0,1]$ across all layers to obtain the final attention importance score $S_{a}^{l}=\frac{\hat{S}_{a}^{l}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}{\max_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}$.
We define a Gaussian prior centered at layer $\mu$ with standard deviation $\sigma$ as $P^{l}=\exp\left(-\frac{(l-\mu)^{2}}{2\sigma^{2}}\right)$. The final selection score for each layer $l$ is computed as a weighted combination of the attention importance score and the Gaussian prior:
$S^{l}=\alpha S_{a}^{l}+(1-\alpha)P^{l},$
where $\alpha\in[0,1]$ is a hyperparameter that balances the two components. We then select the top $M$ layers with the highest selection scores $S^{l}$ to form the subset $\mathcal{S}$ for communication.
For each model pair and dataset, the top $M$ layers are selected based on the selection scores computed from a calibration set. The selected layers are then fixed and used for all samples in the test set. We found that a calibration set as small as a single sample is sufficient to obtain a robust selection that generalizes well to the entire test set, as shown in the experiments in Appendix G.
3.3 Complexity Analysis
We analyze the computational complexity of our KVComm framework compared to baseline methods. Compared to the NLD (Du et al., 2023) method, our method does not require multiple decoding steps for $\mathcal{M}_{s}$, which significantly reduces the computation cost. When the number of tokens generated during debate (Du et al., 2023) is large, the computation margin of our method over NLD is on the order of $O(L(T_{s}+T_{r}+|Q|)^{2}d),$ where $T_{s}$ and $T_{r}$ are the number of tokens generated by $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$ in the debate, respectively, and $|Q|$ and $d$ are the number of tokens in the query and the hidden dimension of the model, respectively. Compared to the Skyline (Section 4.1) method, our method also reduces the computation cost, especially when $M$ is small. The computation margin of our method over Skyline is on the order of O(|C|d(L(2|Q|+T)M(|Q|+T)))O(|C|d(L(2|Q|+T)-M(|Q|+T))), where $|C|$ is the number of tokens in the context, and $|T|$ is the number of tokens generated by $\mathcal{M}_{r}$.
4 Experiments
4.1 Experimental Setup
Datasets
We evaluate KVComm on a diverse set of contextual reasoning tasks. Following Ramesh & Li (2025), we synthetically generate two datasets, Countries, which asks questions about countries based on landmark information, and Tipsheets, which requires investment decisions from financial tips. Examples of these two datasets are shown in Table 3 in Section B.1. Moreover, we select six benchmarks, including HotpotQA (Yang et al., 2018), QASPER (Dasigi et al., 2021), MuSiQuest (Trivedi et al., 2022), two subsets of LongBench (Bai et al., 2024)(MultiFieldQA-en and 2WikiMQA), and TMATH (Qi et al., 2025). The last dataset is a mathematical problem-solving dataset that contains hints as context. We use ROUGE-L Recall as the evaluation metric for the last dataset, and F1 score for all other datasets. Statistics are summarized in Table 4 in Section B.1.
Models
We conduct experiments on eight different model pairs, shown in Table 5 in Section B.3. The model pairs include two instances of the same LLM and two models that are fine-tuned versions of the same base LLM. These models cover different families, including LLaMA (Dubey et al., 2024), Qwen (Qwen et al., 2024), and Falcon (Almazrouei et al., 2023).
Compared Methods
We compare KVComm with several representative approaches: Baseline (no communication between $\mathcal{M}_{r}$ and $\mathcal{M}_{s}$), Skyline (concatenating context $C$ and query $Q$ as an upper bound), Natural Language Debate (NLD) (Du et al., 2023), CIPHER (Pham et al., 2023), and AC (Ramesh & Li, 2025). Detailed descriptions for these methods are provided in Section B.4. Implementation details are provided in Section B.2.
Table 1: Communication results of different methods. Best results are bolded, second best underlined (excluding Baseline and Skyline). We report the results with $\mathcal{M}_{r}$ for Baseline and Skyline for fairness. KVComm (0.3/0.5/0.7) denotes selecting 30%/50%/70% of layers’ KV pairs for communication, i.e., $M=\lceil 0.3L\rceil$, $M=\lceil 0.5L\rceil$, $M=\lceil 0.7L\rceil$.
Method Countries Tipsheets HotpotQA QASPER MuSiQuest
MultiField
-QA-en
2WikiM
-QA
TMATH
$\mathcal{M}_{s}$: huihui-ai/Llama-3.2-3B-Instruct-abliterated; $\mathcal{M}_{r}$: suayptalha/DeepSeek-R1-Distill-Llama-3B
Baseline $0.05$ $0.32$ $0.23$ $0.05$ $0.02$ $0.11$ $0.27$ $0.34$
Skyline $0.57$ $0.91$ $0.73$ $0.25$ $0.51$ $0.47$ $0.40$ $0.36$
NLD $0.03$ 0.73 $0.18$ $0.05$ $0.03$ $0.13$ $0.05$ $0.30$
CIPHER $0.00$ $0.40$ $0.09$ $0.04$ $0.00$ $0.08$ $0.08$ $0.30$
AC (mean) $0.03$ $0.45$ $0.25$ $0.05$ $0.02$ $0.13$ $0.23$ 0.35
AC (replace) $0.00$ $0.49$ $0.05$ $0.01$ $0.01$ $0.12$ $0.03$ 0.34
AC (sum) $0.02$ $0.46$ $0.23$ $0.05$ $0.01$ $0.13$ $0.24$ 0.34
KVComm (0.3) 0.46 $0.45$ $0.46$ $0.09$ $0.28$ $0.15$ $0.28$ 0.35
KVComm (0.5) 0.57 0.81 0.57 0.27 0.32 0.51 0.36 0.35
KVComm (0.7) 0.57 0.81 0.65 0.29 0.36 0.47 0.37 0.35
$\mathcal{M}_{s}$: Orion-zhen/Qwen2.5-7B-Instruct-Uncensored; $\mathcal{M}_{r}$: bespokelabs/Bespoke-Stratos-7B
Baseline $0.01$ $0.36$ $0.13$ $0.05$ $0.03$ $0.08$ $0.09$ $0.35$
Skyline $0.51$ $0.97$ $0.53$ $0.10$ $0.25$ $0.40$ $0.09$ $0.35$
NLD $0.02$ $0.73$ $0.00$ $0.00$ $0.00$ $0.01$ $0.01$ 0.34
CIPHER $0.01$ $0.59$ $0.00$ $0.00$ $0.00$ $0.01$ $0.02$ 0.33
AC (mean) $0.00$ $0.00$ $0.03$ $0.00$ $0.00$ $0.08$ $0.01$ $0.01$
AC (replace) $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$
AC (sum) $0.00$ $0.00$ $0.02$ $0.00$ $0.00$ $0.07$ $0.04$ $0.03$
KVComm (0.3) $0.04$ $0.26$ $0.02$ $0.01$ $0.01$ $0.09$ $0.08$ $0.31$
KVComm (0.5) 0.19 0.88 0.28 0.07 0.12 0.26 0.10 0.33
KVComm (0.7) 0.41 0.89 0.41 0.21 0.25 0.29 0.15 0.34
$\mathcal{M}_{s}$: ehristoforu/falcon3-ultraset; $\mathcal{M}_{r}$: huihui-ai/Falcon3-7B-Instruct-abliterated
Baseline $0.08$ $0.36$ $0.21$ $0.06$ $0.04$ $0.09$ $0.23$ $0.31$
Skyline $0.56$ $0.95$ $0.76$ $0.32$ $0.56$ $0.51$ $0.45$ $0.37$
NLD 0.46 $0.84$ $0.46$ $0.04$ $0.21$ $0.14$ $0.26$ $0.13$
CIPHER $0.31$ $0.18$ $0.18$ $0.01$ $0.05$ $0.06$ $0.26$ $0.11$
AC (mean) $0.01$ $0.46$ $0.25$ $0.06$ $0.04$ $0.09$ $0.23$ $0.31$
AC (replace) $0.00$ $0.49$ $0.12$ $0.00$ $0.01$ $0.13$ $0.17$ $0.31$
AC (sum) $0.01$ $0.46$ $0.25$ $0.06$ $0.03$ $0.10$ $0.24$ $0.31$
KVComm (0.3) 0.46 $0.69$ 0.59 $0.19$ $0.40$ $0.35$ $0.29$ $0.32$
KVComm (0.5) 0.40 0.92 0.63 0.25 0.44 0.45 0.34 0.35
KVComm (0.7) $0.19$ 0.96 $0.55$ 0.26 0.42 0.51 0.31 0.36
4.2 Communication Results
Table 1 reports results on three model pairs fine-tuned from the same base LLM. The results on other model pairs are provided in LABEL:tab:more_communication_results in Appendix E, which show similar trends. We observe that KVComm consistently outperforms all baseline communication methods across datasets and model pairs. AC can outperform the Baseline method on some datasets, but they are still significantly worse than KVComm and Skyline, as hidden states of $\mathcal{M}_{r}$ are corrupted during communication. NLD and CIPHER methods perform poorly on most datasets, as they are designed for improving the answer quality through debate rather than communication. $\mathcal{M}_{s}$ cannot effectively convey useful information to $\mathcal{M}_{r}$ through debate, as $\mathcal{M}_{s}$ has no information about the query $Q$. Our KVComm framework can achieve comparable performance than the Skyline method when selecting 70% of layers’ KV pairs for communication, demonstrating the effectiveness of our selection strategy. Even when selecting only 30% of layers’ KV pairs, KVComm can still outperform most baseline communication methods on many datasets, showing its potential for efficient communication with minimal overhead.
Note that KVComm can outperform Skyline on some datasets. We attribute this to two factors: (1) $\mathcal{M}_{s}$ may complement $\mathcal{M}_{r}$ with stronger capabilities in certain aspects, and (2) selective KV sharing provides a regularization effect, which helps $\mathcal{M}_{r}$ to focus on the most relevant information and avoid wasting its capacity on less important signals. This also explains why using fewer layers can sometimes yield better performance than using more.
Also note that the performance gain of KVComm is not substantial on TMATH. We attribute this to that pretraining gives LLMs solid capabilities in mathematical reasoning, which may not dramatically benefit from additional context or hints. Moreover, AC performs relatively well on this dataset, which we consider is because the hints contain information about questions, so even if the last token’s hidden states are corrupted, it can still generate some useful information.
4.3 Benefit of Selective KV Over One Contiguous Chunk
DroidSpeak (Liu et al., 2024b) chooses to use one contiguous chunk of context for communication between LLMs. Despite different problem settings, we evaluate KVComm by replacing the selection strategy with two hyperparameters, which are two layer indices $\textrm{layer}_{\textrm{from}}$ and $\textrm{layer}_{\textrm{to}}$, then all layers between $\textrm{layer}_{\textrm{from}}$ and $\textrm{layer}_{\textrm{to}}$ are selected for communication. This is equivalent to using one contiguous chunk of context for communication. We vary them to select different chunks of layers.
Figure 6 shows that using a single contiguous chunk for communication yields good performance only in a small region of the hyperparameter space, making it tricky to find the right hyperparameters. In contrast, the scatter and curve plots in Figure 6 demonstrate that KVComm consistently achieves the best or even outperforms the best contiguous chunk setting for the same number of layers. Line plots in Figure 6 show that contiguous chunks are most effective when taken from intermediate layers, consistent with hypothesis H1 in Section 3.2. All results are on HotpotQA with the Llama-3.1-8B pair, with more in Appendix I.
4.4 Ablation Study on Selection Strategy
Table 2 compares KVComm with random selection. We find that KVComm consistently outperforms random selection across different datasets and selection ratios. When the ratio is high (i.e., 0.7), the performance gap between our selection strategy and random selection becomes smaller, as more layers are selected and the impact of the selection strategy is reduced. However, when the ratio is low (i.e., 0.3), our selection strategy significantly outperforms random selection, demonstrating its effectiveness in selecting the most informative layers for communication. Comparison results on other model pairs are provided in LABEL:tab:more_comparison_with_random in Appendix F, which show similar trends.
Table 2: Comparison with random selection. Best results for each selection ratio are bolded.
Method Countries Tipsheets HotpotQA QASPER MuSiQuest
MultiField
-QA-en
2WikiM
-QA
TMATH
$\mathcal{M}_{s}$: huihui-ai/Llama-3.2-3B-Instruct-abliterated; $\mathcal{M}_{r}$: suayptalha/DeepSeek-R1-Distill-Llama-3B
Random (0.3) $0.05$ $0.32$ $0.18$ $0.07$ $0.01$ $0.06$ $0.17$ $0.33$
KVComm (0.3) 0.46 0.45 0.46 0.09 0.28 0.15 0.28 0.35
Random (0.5) $0.26$ $0.44$ $0.37$ $0.08$ $0.10$ $0.09$ $0.21$ $0.34$
KVComm (0.5) 0.57 0.81 0.57 0.27 0.32 0.51 0.36 0.35
Random (0.7) 0.57 0.82 $0.62$ $0.20$ $0.34$ $0.30$ $0.28$ 0.35
KVComm (0.7) 0.57 $0.81$ 0.65 0.29 0.36 0.47 0.37 0.35
Figure 4: Effective communication with limited hyperparameters.
Figure 5: KVComm achieves nearly the best or even outperforms contig. chunks.
Figure 6: Chunks in intermediate layers achieve the most effective communication.
4.5 Attention Distribution Analysis
We validate hypothesis H2 in Section 3.2 by selecting layers with different attention importance scores for communication. We select $9$ layers with different levels of attention importance scores, and test the communication performance with Llama-3.2-3B model. The results are shown in Figure 8. We can find that selecting layers with higher scores can achieve better performance, while selecting layers with lower scores can diminish the performance. This validates hypothesis H2 that layers with higher attention importance scores are more effective for communication.
4.6 System Efficiency
Mathematically, we have shown in Section 3.3 that KVComm can reduce the computation cost compared to Skyline. We validate this through experiments on the Llama-3.2-3B model pair with Tipsheets and MultiFieldQA-en datasets. We report the relative FLOPs of KVComm and Skyline over AC in Figure 8. NLD and CIPHER are not included since they require multiple decoding steps for $\mathcal{M}_{s}$, which makes the computation cost significantly higher than AC. We can find that KVComm has a significant computation advantage over Skyline, especially when selecting fewer layers for communication. This demonstrates the efficiency of our KVComm framework in enabling effective communication with reduced computational overhead by 2.5x to 6x.
Figure 7: Better communication performance with higher attention level.
Figure 8: KVComm requires less computation compared to Skyline.
5 Related Work
LLM Inference Acceleration
Lots of work has focused on accelerating LLM inference. Computation-level methods such as FlashAttention (Dao et al., 2022) and Memory-Efficient Attention (Rabe & Staats, 2021) reduce memory and speed up attention; system-level methods such as vLLM (Kwon et al., 2023) and DeepSpeed-Inference (Aminabadi et al., 2022) improve overall throughput and latency; and model-level methods such as quantization (Lin et al., 2024) and pruning (Ma et al., 2023) reduce model size and complexity. These approaches are orthogonal to ours and can be combined with KVComm to further improve efficiency.
Closest to our work are methods that reuse computation across decoding steps or requests. Gao et al. (2024) introduces a hierarchical KV caching system for all requests; Gim et al. (2024) reuses prompt KV caches across queries by decomposing inputs; Liu et al. (2024c) compresses KV caches into compact bitstreams; and Yao et al. (2025) combines multiple chunks’ KV caches by selectively recomputing a few tokens. In contrast, our work targets communication across different LLMs, which is more challenging due to parameter differences. Moreover, while prior methods reuse KV caches uniformly across layers, we enable selective sharing of KV caches from different layers, further improving efficiency. We do not compare with these works since they are orthogonal to ours.
DroidSpeak (Liu et al., 2024b) also explores KV cache reuse across LLMs, but with key differences. Their goal is to accelerate inference for multiple queries with the same prefix, while we focus on efficient communication between two LLMs on contextual tasks. They reuse a single contiguous chunk of layers and recompute the rest, whereas our strategy flexibly selects non-contiguous layers based on attention importance and a Gaussian prior. Moreover, their method incurs extra overhead from recomputation, while ours directly integrates the selected KV pairs into $\mathcal{M}_{r}$ without recomputation. Despite the different problem settings, we compare their contiguous-chunk strategy with ours in Section 4.3, showing the advantages of our approach.
Inter-LLM Communication
Communication between multiple LLMs has been explored in several recent works. Most works focus on using natural language as the medium of communication. For example, Du et al. (2023) proposed a natural language debate framework where LLMs iteratively critique each other’s answers in natural language to improve the final answer. Liang et al. (2023) followed a similar idea but introduced a judge model to manage the debate process.
CIPHER (Pham et al., 2023) proposed using embedding space as the medium of communication. They pass the weighted average of the token embeddings from one LLM to another. Moreover, AC (Ramesh & Li, 2025) proposed to use the last token’s hidden state as the medium of communication. They replace the last token’s hidden state of the receiver model with that of the sender model. Instead, we propose to use the KV pairs as the medium, which can preserve more information than just using the last token’s hidden state. We also propose a more effective selection strategy for choosing which KV pairs to share, which can further improve efficiency.
KV Cache Optimization
Several works have explored optimizing KV caches for a single LLM by (1) compressing the KV caches to reduce memory usage (Ge et al., 2023; Liu et al., 2024a) or (2) managing the KV caches (offloading) to improve the inference speed (Lee et al., 2024; Xiong et al., 2024). As our work focuses on layer-wise selection of KV caches for communication between two LLMs, these methods are orthogonal and can be combined with our method.
6 Conclusion
In this work, we identified the potential of using KV pairs as an effective medium for communication between two LLMs. We proposed a novel KVComm framework that enables efficient communication by selectively sharing KV pairs between LLM models. We designed a selection strategy based on attention importance scores and a Gaussian prior to select the most relevant layers. Extensive experiments on diverse datasets and model pairs demonstrated that KVComm can achieve comparable or even superior performance to the Skyline upper bound and other methods, while reducing communication costs by up to 3x. We highlight the generalization ability of our selection strategy, which can be effectively calibrated with only a single sample. Our work opens up new possibilities for efficient inter-LLM communication and paves the way for future research in this direction.
Appendix A The Use of Large Language Models (LLMs)
Large language models, including ChatGPT, were employed to provide assistance in improving the clarity, coherence, and fluency of the manuscript. These tools were used solely for language refinement, and all scientific content and interpretations remain the responsibility of the authors.
Appendix B Experimental Setup
B.1 Dataset
We provide sample prompts and expected answers for the Countries and Tipsheets datasets in Table 3, which are inspired by Ramesh & Li (2025). We also provide the statistics of all datasets used in our experiments in Table 4. HotpotQA, QASPER, MuSiQuest, and TMATH datasets are randomly sampled from their original datasets to reduce the evaluation cost.
Table 3: Sample prompts and expected answers for Countries and Tipsheets datasets inspired by Ramesh & Li (2025).
Dataset Role Content
Countries $C$ Uma is at the Mahaffie House.
$Q$ Which country is Uma located in?
Answer United States
Tipsheets $C$ Atlas LLC is under pressure amid softer trends; EPS -17%; won a sizable customer contract but faces a lawsuit. Sable LLC shows clear momentum and improving execution; authorized a buyback but reported a cyber incident. Trace LLC looks balanced with a mixed near-term setup.
$Q$ You must invest in exactly one company from Atlas LLC, Sable LLC, Trace LLC. Which do you choose?
Answer Sable LLC
Table 4: Statistics of the datasets in our experiments.
Dataset Size
Countries 200
Tipsheets 500
HotpotQA (Yang et al., 2018) 500
QASPER (Dasigi et al., 2021) 500
MuSiQuest (Trivedi et al., 2022) 500
MultiFieldQA-en (Bai et al., 2024) 150
2WikiMQA (Bai et al., 2024) 200
TMATH (Qi et al., 2025) 300
B.2 Implementation Details
[…]
B.3 Model Pairs
We conduct experiments on eight different model pairs, shown in Table 5. The first four pairs consist of the same LLMs, while the last four pairs consist of models that are fine-tuned on the same base LLM.
Table 5: Model pairs in the evaluation. $\mathcal{M}_{s}$ is the sender model, and $\mathcal{M}_{r}$ is the receiver model.
Index $\mathcal{M}_{s}$ $\mathcal{M}_{r}$ Note
1 meta-llama/Llama-3.1-8B-Instruct meta-llama/Llama-3.1-8B-Instruct Same model
2 meta-llama/Llama-3.2-3B-Instruct meta-llama/Llama-3.2-3B-Instruct Same model
3 Qwen/Qwen2.5-7B-Instruct Qwen/Qwen2.5-7B-Instruct Same model
4 tiiuae/Falcon3-7B-Instruct tiiuae/Falcon3-7B-Instruct Same model
5 yuvraj17/EvolCodeLlama-3.1-8B-Instruct Team-ACE/ToolACE-2-Llama-3.1-8B Fine-tuned on 1
6 huihui-ai/Llama-3.2-3B-Instruct-abliterated suayptalha/DeepSeek-R1-Distill-Llama-3B Fine-tuned on 2
7 Orion-zhen/Qwen2.5-7B-Instruct-Uncensored bespokelabs/Bespoke-Stratos-7B Fine-tuned on 3
8 ehristoforu/falcon3-ultraset huihui-ai/Falcon3-7B-Instruct-abliterated Fine-tuned on 4
B.4 Compared Method Descriptions
We compare our proposed KVComm framework with the following methods:
  • Baseline: $\mathcal{M}_{r}$ processes the query $Q$ without any communication from $\mathcal{M}_{s}$.
  • Skyline: $\mathcal{M}_{r}$ directly processes the concatenation of the context $C$ and query $Q$. This serves as an upper bound for performance.
  • Natural Language Debate (NLD) (Du et al., 2023): Each model generates an initial answer, and then they iteratively critique each other’s answers in natural language for a fixed number of rounds. Finally, one model produces the final answer based on the entire debate history. Compared to the original setting, we explicitly tell $\mathcal{M}_{s}$ that it has to summarize the context $C$ in its initial answer. We set the number of debate rounds to 2.
  • CIPHER (Pham et al., 2023): Similar to NLD, but instead of communicating in natural language, the models communicate by passing the weighted average of the token embeddings from one LLM to another. We use the same prompt as NLD, and set the number of debate rounds to 2.
  • AC (Ramesh & Li, 2025): Communicate with the last token’s hidden state. Replace the last token’s hidden state of $\mathcal{M}_{r}$ with that of $\mathcal{M}_{s}$. We also test with mean and sum operations.
Appendix C Token Importance at Different Positions
We conduct the same experiment as in Section 2.2.1 on other datasets and models to investigate the effect of tokens at different positions in the sequence on the model’s output. We report the results on MMLU Social Science, MMLU STEM, and MMLU Humanities using Llama-3.1-8B and Llama-3.2-3B models in Figure 9. We can see that the last token’s hidden state plays the most critical role in the latter layers, which is consistent with the observation in Section 2.2.1.
(a) Llama-3.2-3B on MMLU Social Science
(b) Llama-3.1-8B on MMLU STEM
(c) Llama-3.2-3B on MMLU STEM
(d) Llama-3.1-8B on MMLU Humanities
(e) Llama-3.2-3B on MMLU Humanities
Figure 9: Effect of removing or retaining a token’s hidden state across different positions on MMLU Social Science, MMLU STEM, and MMLU Humanities accuracy using Llama-3.1-8B and Llama-3.2-3B models.
Appendix D Utilizing All Tokens
We conduct the same experiment as in Section 2.2.2 on Countries, Tipsheets, and HotpotQA datasets using Llama-3.1-8B, Llama-3.2-3B, and Qwen2.5-7B models. The results are shown in Figure 10. We can see the results are consistent with the observation in Section 2.2.2.
(a) Countries
(b) Tipsheets
(c) HotpotQA
Figure 10: Performance heatmap of prepending the hidden states from certain layers of $\mathcal{M}_{s}$ to certain layers of $\mathcal{M}_{r}$ on Countries, Tipsheets, and HotpotQA.
Appendix E More Communication Results
We provide more communication results on different model pairs in LABEL:tab:more_communication_results, which show similar trends as in Section 4.2.
Table 6: More communication results of different methods. Best results are bolded, second best underlined (excluding Baseline and Skyline). We report $\mathcal{M}_{r}$ for Baseline and Skyline for fairness. KVComm (0.3/0.5/0.7) denotes selecting 30%/50%/70% of layers’ KV pairs for communication, i.e., $M=\lceil 0.3L\rceil$, $M=\lceil 0.5L\rceil$, $M=\lceil 0.7L\rceil$.
Method Countries Tipsheets HotpotQA QASPER MuSiQuest
MultiField
-QA-en
2WikiM
-QA
TMATH
$\mathcal{M}_{s}$: meta-llama/Llama-3.1-8B-Instruct; $\mathcal{M}_{r}$: meta-llama/Llama-3.1-8B-Instruct
Baseline $0.00$ $0.05$ $0.19$ $0.02$ $0.01$ $0.07$ $0.06$ $0.35$
Skyline $0.62$ $0.92$ $0.74$ $0.35$ $0.54$ $0.56$ $0.52$ $0.36$
NLD $0.00$ $0.85$ $0.06$ $0.02$ $0.00$ $0.05$ $0.03$ $0.36$
CIPHER $0.03$ $0.82$ $0.10$ $0.01$ $0.01$ $0.10$ $0.02$ $0.36$
AC (mean) $0.00$ $0.12$ $0.19$ $0.02$ $0.01$ $0.08$ $0.03$ $0.35$
AC (replace) $0.00$ $0.36$ $0.15$ $0.02$ $0.01$ $0.07$ $0.05$ $0.35$
AC (sum) $0.00$ $0.09$ $0.20$ $0.02$ $0.01$ $0.09$ $0.04$ $0.35$
KVComm (0.3) 0.51 $0.93$ $0.33$ 0.07 $0.11$ $0.21$ $0.29$ 0.37
KVComm (0.5) 0.62 0.95 0.60 0.29 0.34 0.50 0.37 0.37
KVComm (0.7) 0.62 0.96 0.69 0.29 0.39 0.53 0.38 0.38
$\mathcal{M}_{s}$: meta-llama/Llama-3.2-3B-Instruct; $\mathcal{M}_{r}$: meta-llama/Llama-3.2-3B-Instruct
Baseline $0.02$ $0.01$ $0.16$ $0.00$ $0.02$ $0.10$ $0.09$ $0.35$
Skyline $0.56$ $0.87$ $0.72$ $0.23$ $0.45$ $0.45$ $0.37$ $0.38$
NLD $0.00$ $0.14$ $0.06$ $0.01$ $0.00$ $0.03$ $0.00$ $0.29$
CIPHER $0.02$ $0.45$ $0.07$ $0.02$ $0.01$ $0.02$ $0.01$ $0.31$
AC (mean) $0.00$ $0.07$ $0.18$ $0.01$ $0.02$ $0.09$ $0.06$ $0.35$
AC (replace) $0.01$ $0.37$ $0.13$ $0.01$ $0.02$ $0.06$ $0.03$ $0.34$
AC (sum) $0.00$ $0.34$ $0.20$ $0.02$ $0.02$ $0.10$ $0.07$ $0.34$
KVComm (0.3) $0.51$ $0.48$ $0.47$ $0.10$ $0.20$ $0.17$ $0.28$ $0.35$
KVComm (0.5) 0.55 0.79 0.58 0.24 0.27 0.47 0.35 0.36
KVComm (0.7) 0.57 0.80 0.65 0.27 0.29 0.48 0.31 0.37
$\mathcal{M}_{s}$: Qwen/Qwen2.5-7B-Instruct; $\mathcal{M}_{r}$: Qwen/Qwen2.5-7B-Instruct
Baseline $0.00$ $0.32$ $0.19$ $0.05$ $0.03$ $0.06$ $0.17$ $0.32$
Skyline $0.54$ $0.97$ $0.68$ $0.30$ $0.48$ $0.49$ $0.45$ $0.33$
NLD $0.00$ $0.88$ $0.00$ $0.00$ $0.00$ $0.01$ $0.00$ $0.30$
CIPHER $0.00$ $0.89$ $0.01$ $0.00$ $0.00$ $0.03$ $0.00$ $0.31$
AC (mean) $0.00$ $0.37$ $0.15$ $0.01$ $0.02$ $0.10$ $0.20$ 0.33
AC (replace) $0.00$ $0.35$ $0.02$ $0.00$ $0.00$ $0.10$ $0.09$ 0.32
AC (sum) $0.00$ $0.41$ $0.14$ $0.02$ $0.02$ $0.08$ $0.17$ 0.32
KVComm (0.3) $0.04$ $0.31$ $0.06$ $0.02$ $0.01$ $0.19$ $0.19$ 0.32
KVComm (0.5) 0.57 0.92 0.49 0.18 0.20 0.40 0.25 0.32
KVComm (0.7) 0.56 0.98 0.72 0.29 0.48 0.45 0.35 0.33
$\mathcal{M}_{s}$: tiiuae/Falcon3-7B-Instruct; $\mathcal{M}_{r}$: tiiuae/Falcon3-7B-Instruct
Baseline $0.06$ $0.33$ $0.19$ $0.04$ $0.04$ $0.09$ $0.21$ $0.31$
Skyline $0.57$ $0.95$ $0.70$ $0.24$ $0.50$ $0.49$ $0.48$ $0.35$
NLD $0.39$ $0.79$ $0.33$ $0.02$ $0.11$ $0.13$ $0.27$ $0.18$
CIPHER $0.45$ $0.68$ $0.24$ $0.00$ $0.07$ $0.08$ $0.24$ $0.19$
AC (mean) $0.03$ $0.51$ $0.22$ $0.04$ $0.04$ $0.09$ $0.22$ 0.32
AC (replace) $0.00$ $0.57$ $0.09$ $0.00$ $0.02$ $0.12$ $0.14$ 0.31
AC (sum) $0.04$ $0.51$ $0.22$ $0.04$ $0.03$ $0.09$ $0.22$ 0.32
KVComm (0.3) $0.06$ $0.67$ $0.41$ $0.12$ $0.22$ $0.41$ $0.23$ 0.32
KVComm (0.5) $0.16$ $0.94$ $0.52$ 0.22 0.33 0.47 0.33 0.32
KVComm (0.7) 0.23 0.96 0.54 0.22 $0.32$ 0.47 $0.29$ 0.32
$\mathcal{M}_{s}$: yuvraj17/EvolCodeLlama-3.1-8B-Instruct; $\mathcal{M}_{r}$: Team-ACE/ToolACE-2-Llama-3.1-8B
Baseline $0.00$ $0.07$ $0.04$ $0.00$ $0.01$ $0.08$ $0.01$ $0.34$
Skyline $0.24$ $0.95$ $0.37$ $0.17$ $0.15$ $0.51$ $0.25$ $0.39$
NLD $0.00$ $0.82$ $0.05$ $0.03$ $0.01$ $0.10$ $0.02$ $0.29$
CIPHER $0.00$ $0.86$ $0.05$ $0.01$ $0.02$ $0.09$ $0.01$ $0.31$
AC (mean) $0.00$ $0.31$ $0.03$ $0.00$ $0.01$ $0.11$ $0.01$ $0.34$
AC (replace) $0.00$ $0.30$ $0.05$ $0.00$ $0.01$ $0.10$ $0.02$ $0.33$
AC (sum) $0.00$ $0.27$ $0.04$ $0.00$ $0.01$ $0.09$ $0.01$ $0.34$
KVComm (0.3) $0.12$ $0.95$ $0.12$ $0.05$ $0.04$ $0.26$ $0.19$ 0.36
KVComm (0.5) 0.55 0.98 0.38 0.15 0.14 0.43 0.28 0.38
KVComm (0.7) 0.53 0.97 0.51 0.22 0.25 0.49 0.33 0.38
Appendix F Ablation Study on Selection Strategy
We conduct more ablation studies on the selection strategy by comparing with random selection and selection based on only attention importance scores. The results are shown in LABEL:tab:more_comparison_with_random, which show similar trends as in Section 4.4.
Table 7: More comparison results with random selection. Best results for each selection ratio are bolded.
Method Countries Tipsheets HotpotQA QASPER MuSiQuest
MultiField
-QA-en
2WikiM
-QA
TMATH
$\mathcal{M}_{s}$: meta-llama/Llama-3.1-8B-Instruct; $\mathcal{M}_{r}$: meta-llama/Llama-3.1-8B-Instruct
Random (0.3) $0.02$ $0.35$ $0.24$ 0.07 $0.04$ $0.07$ $0.12$ $0.35$
KVComm (0.3) 0.51 0.93 0.33 0.07 0.11 0.21 0.29 0.37
Random (0.5) $0.49$ $0.76$ $0.58$ $0.15$ $0.29$ $0.29$ $0.27$ $0.36$
KVComm (0.5) 0.62 0.95 0.60 0.29 0.34 0.50 0.37 0.37
Random (0.7) 0.63 $0.88$ 0.76 0.32 0.49 $0.52$ $0.34$ $0.37$
KVComm (0.7) $0.62$ 0.96 $0.69$ $0.29$ $0.39$ 0.53 0.38 0.38
$\mathcal{M}_{s}$: Orion-zhen/Qwen2.5-7B-Instruct-Uncensored; $\mathcal{M}_{r}$: bespokelabs/Bespoke-Stratos-7B
Random (0.3) $0.00$ $0.09$ $0.00$ $0.00$ $0.00$ $0.06$ $0.01$ 0.31
KVComm (0.3) 0.04 0.26 0.02 0.01 0.01 0.09 0.08 0.31
Random (0.5) $0.12$ $0.32$ $0.06$ $0.00$ $0.03$ $0.15$ $0.04$ 0.33
KVComm (0.5) 0.19 0.88 0.28 0.07 0.12 0.26 0.10 0.33
Random (0.7) $0.16$ $0.76$ $0.14$ $0.03$ $0.02$ $0.20$ $0.04$ 0.34
KVComm (0.7) 0.41 0.89 0.41 0.21 0.25 0.29 0.15 0.34
$\mathcal{M}_{s}$: ehristoforu/falcon3-ultraset; $\mathcal{M}_{r}$: huihui-ai/Falcon3-7B-Instruct-abliterated
Random (0.3) $0.35$ $0.36$ $0.23$ $0.06$ $0.07$ $0.14$ $0.24$ 0.31
KVComm (0.3) 0.46 0.69 0.59 0.19 0.40 0.35 0.29 0.32
Random (0.5) $0.23$ $0.42$ $0.27$ $0.09$ $0.08$ $0.15$ $0.28$ $0.31$
KVComm (0.5) 0.40 0.92 0.63 0.25 0.44 0.45 0.34 0.35
Random (0.7) $0.18$ $0.94$ $0.51$ $0.23$ $0.35$ $0.47$ $0.30$ $0.34$
KVComm (0.7) 0.19 0.96 0.55 0.26 0.42 0.51 0.31 0.36
$\mathcal{M}_{s}$: meta-llama/Llama-3.2-3B-Instruct; $\mathcal{M}_{r}$: meta-llama/Llama-3.2-3B-Instruct
Random (0.3) $0.02$ $0.29$ $0.11$ $0.06$ $0.02$ $0.07$ $0.16$ $0.34$
KVComm (0.3) 0.51 0.48 0.47 0.10 0.20 0.17 0.28 0.35
Random (0.5) $0.28$ $0.44$ $0.30$ $0.06$ $0.06$ $0.06$ $0.19$ $0.35$
KVComm (0.5) 0.55 0.79 0.58 0.24 0.27 0.47 0.35 0.36
Random (0.7) $0.54$ 0.81 $0.62$ $0.21$ 0.30 $0.30$ $0.26$ $0.36$
KVComm (0.7) 0.57 $0.80$ 0.65 0.27 $0.29$ 0.48 0.31 0.37
$\mathcal{M}_{s}$: Qwen/Qwen2.5-7B-Instruct; $\mathcal{M}_{r}$: Qwen/Qwen2.5-7B-Instruct
Random (0.3) $0.00$ $0.34$ $0.05$ $0.00$ $0.00$ $0.08$ $0.10$ 0.30
KVComm (0.3) 0.04 0.31 0.06 0.02 0.01 0.19 0.19 0.32
Random (0.5) $0.00$ $0.32$ $0.10$ $0.02$ $0.02$ $0.10$ $0.16$ 0.32
KVComm (0.5) 0.57 0.92 0.49 0.18 0.20 0.40 0.25 0.32
Random (0.7) $0.41$ $0.71$ $0.28$ $0.04$ $0.04$ $0.21$ $0.17$ $0.32$
KVComm (0.7) 0.56 0.98 0.72 0.29 0.48 0.45 0.35 0.33
$\mathcal{M}_{s}$: tiiuae/Falcon3-7B-Instruct; $\mathcal{M}_{r}$: tiiuae/Falcon3-7B-Instruct
Random (0.3) $0.01$ $0.35$ $0.18$ $0.04$ $0.03$ $0.12$ $0.21$ $0.30$
KVComm (0.3) 0.06 0.67 0.41 0.12 0.22 0.41 0.23 0.32
Random (0.5) $0.04$ $0.41$ $0.24$ $0.03$ $0.05$ $0.16$ $0.24$ $0.31$
KVComm (0.5) 0.16 0.94 0.52 0.22 0.33 0.47 0.33 0.32
Random (0.7) $0.19$ $0.95$ $0.51$ $0.20$ $0.29$ $0.42$ $0.26$ 0.32
KVComm (0.7) 0.23 0.96 0.54 0.22 0.32 0.47 0.29 0.32
$\mathcal{M}_{s}$: yuvraj17/EvolCodeLlama-3.1-8B-Instruct; $\mathcal{M}_{r}$: Team-ACE/ToolACE-2-Llama-3.1-8B
Random (0.3) $0.00$ $0.34$ $0.06$ $0.00$ $0.01$ $0.13$ $0.03$ $0.34$
KVComm (0.3) 0.12 0.95 0.12 0.05 0.04 0.26 0.19 0.36
Random (0.5) $0.03$ $0.79$ $0.29$ $0.06$ $0.09$ $0.32$ $0.16$ $0.35$
KVComm (0.5) 0.55 0.98 0.38 0.15 0.14 0.43 0.28 0.38
Random (0.7) $0.37$ $0.85$ 0.59 $0.21$ 0.27 $0.47$ 0.33 $0.36$
KVComm (0.7) 0.53 0.97 $0.51$ 0.22 $0.25$ 0.49 0.33 0.38
Appendix G Calibration Set Size
Figure 11: Effect of calibration set size. Calibration set size does not significantly affect the test performance.
We investigate how many samples are needed in the calibration set so that the selection strategy can generalize well to the test set. If a smaller calibration set can achieve good performance on the test set, it would be more practical since it would require less cost to obtain the selected layers. We conduct the experiment on Countries, Tipsheets, and HotpotQA datasets using the Llama-3.2-3B model. As the results in Figure 11 show, we can see that using only one sample in the calibration set can already achieve the same performance as using more samples (up to 128 samples). This suggests that our selection strategy can generalize well to the test set even with a very small calibration set. In all other experiments in the paper, we use one sample in the calibration set.
Appendix H Complexity Analysis Details
We compare the computational complexity of our KVComm framework with the Skyline method and the NLD method. Recall that $L$ is the total number of layers in the model, $M$ is the number of selected layers for communication. We use $d$ to denote the hidden dimension of the model, and $|Q|$ and $|C|$ to denote the number of tokens in the query and context, respectively. Suppose $\mathcal{M}_{r}$ would generate $T$ tokens in total, and the number of generated tokens is the same across different methods. For NLD, $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$ would each generate $T_{s}$ and $T_{r}$ tokens for the initial answer, respectively.
Ignoring the embedding, output layers, and other minor components, the computational complexity of prefilling a sequence of length $N$ with a single decoder layer is $O(Nd^{2}+N^{2}d)$, while the complexity of decoding a single token is $O(d^{2}+(N+i)d)$, where $i$ is the index of the generated token. Therefore, the total computational complexity of $\mathcal{M}_{s}$ to process the context $C$ is $O(L(|C|d^{2}+|C|^{2}d))$.
The total computational complexity of KVComm consists of three parts: (1) the complexity of $\mathcal{M}_{s}$ to process the context $C$, which is $O(L(|C|d^{2}+|C|^{2}d))$, (2) the complexity of $\mathcal{M}_{r}$ to process the query $Q$ with the selected $M$ KV pairs from $\mathcal{M}_{s}$, which is $O(L|Q|d^{2}+M(|C|+|Q|)|Q|d+(L-M)|Q|^{2}d)$, and (3) the complexity of $\mathcal{M}_{r}$ to generate $T$ tokens with the selected $M$ KV pairs from $\mathcal{M}_{s}$, which is $O(T(Ld^{2}+M(|C|+|Q|+T)d+(L-M)(|Q|+T)d))$. Therefore, the total computational complexity of KVComm is:
$T(\text{KVComm})=$ $\;O\!\Big(L\left(|C|+|Q|+T\right)d^{2}\Big)$
$+O\!\Big(\big(L\left(|C|^{2}+|Q|^{2}+T^{2}+T|Q|\right)+CM\left(|Q|+T\right)\big)d\Big)$
The computational complexity of Skyline method consists of two parts: (1) the complexity of prefilling the concatenation of the context $C$ and query $Q$, which is $O(L(|C|+|Q|)d^{2}+L(|C|+|Q|)^{2}d)$, and (2) the complexity of decoding $T$ tokens, which is $O(TL(d^{2}+(|C|+|Q|+T)d))$. Therefore, the total computational complexity of the Skyline method is:
$T(\text{Skyline})=$ $\;O\!\Big(L\big(|C|+|Q|+T\big)d^{2}\Big)$
$+O\!\Big(L\Big((|C|+|Q|)^{2}+T\big(|C|+|Q|+T\big)\Big)d\Big)$
The margin of KVComm over Skyline is:
$T(\text{Skyline})-T(\text{KVComm})=$ $\;O\!\Big(|C|d\big(L(2|Q|+T)-M(|Q|+T)\big)\Big)$
For NLD, the total computational complexity consists of three parts: (1) the complexity of $\mathcal{M}_{s}$ to process the context $C$ and generate $T_{s}$ tokens, which is $O(L(|C|d^{2}+|C|^{2}d)+T_{s}L(d^{2}+(|C|+T_{s})d))$, (2) the complexity of $\mathcal{M}_{r}$ to process the query $Q$ and generate $T_{r}$ tokens, which is $O(L(|Q|d^{2}+|Q|^{2}d)+T_{r}L(d^{2}+(|Q|+T_{r})d))$, and (3) the complexity of $\mathcal{M}_{r}$ to process the entire debate history and generate $T$ tokens, which is $O(L((T_{s}+T_{r}+|Q|)d^{2}+(T_{s}+T_{r}+|Q|)^{2}d)+TL(d^{2}+(T_{s}+T_{r}+|Q|+T)d))$. Therefore, the total computational complexity of NLD is:
$T(\text{NLD})=$ $\;O\!\Bigg(L\Big(|C|+2|Q|+2T_{s}+2T_{r}+T\Big)d^{2}\Bigg)$
$+O\!\Bigg(L\Big(|C|^{2}+T_{s}^{2}+|Q|^{2}+T_{r}^{2}+\big(T_{s}+T_{r}+|Q|\big)^{2}$
$\quad\quad+T\big(T_{s}+T_{r}+T+|Q|\big)+T_{s}|C|+T_{r}|Q|\Big)d\Bigg)$
The margin of KVComm over NLD is:
$T(\text{NLD})-T(\text{KVComm})=$ $\;O\!\Big(L\big(2T_{s}+2T_{r}+|Q|\big)d^{2}\Big)$
$+O\!\Bigg(\Bigg(L\Big(T_{s}^{2}+T_{r}^{2}+\big(T_{s}+T_{r}+|Q|\big)^{2}$
$+T_{s}|C|+T_{r}|Q|+T(T_{s}+T_{r})\Big)-CM\big(|Q|+T\big)\Bigg)d\Bigg)$
Appendix I Using One Chunk of Layers
We conduct the same experiment as in Section 4.3 on the HotpotQA dataset using other model pairs in Table 5. The results are shown in Figure 13. We can see that the results are consistent with the observation in Section 4.3.
(a) Llama-3.2-3B-Instruct as both $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$
(b) Qwen2.5-7B-Instruct as both $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$
(c) Qwen2.5-7B-Instruct-Uncensored as $\mathcal{M}_{s}$ and Bespoke-Stratos-7B as $\mathcal{M}_{r}$
(d) EvolCodeLlama-3.1-8B-Instruct as $\mathcal{M}_{s}$ and ToolACE-2-Llama-3.1-8B as $\mathcal{M}_{r}$
(e) Llama-3.2-3B-Instruct-abliterated as $\mathcal{M}_{s}$ and DeepSeek-R1-Distill-Llama-3B as $\mathcal{M}_{r}$
Figure 13: Experiment results of using one chunk of layers for communication on HotpotQA dataset using different model pairs.